# ObjectRelator: Enabling Cross-View Object Relation Understanding in Ego-Centric and Exo-Centric Videos

### v1

@@ 3.2 Key Module: Multimodal Condition Fusion
Our multimodal condition fusion (MCFuse) is mainly presented to break the limitation of the base model that only visual prompt is used, achieving the effective combination of the embedding from both text condition and visual mask condition. There are mainly two considerations for the design of MCFuse, 1) Given the extensive effectiveness of cross-attention in capturing intricate relationships across modalities, it serves as an ideal choice for fusing our multimodal conditions; 2) the model-generated texts can sometimes be unreliable, prioritizing visual information is prudent. This motivates our use of a residual design, with the visual input serving as the primary pathway. Additionally, to eliminate the need for manual tuning of the residual pathway’s weight, we introduce a learnable parameter, $k_{lea}$.
@@
Formally, as shown in Fig. 3, given the ego text emb $E^{ego}_{T}\in 1\times D$ and ego visual emb $E^{ego}_{V}\in N\times D$ as input, the cross-attention is applied to the broadcasted $E^{ego}_{T}$ and $E^{ego}_{V}$, with $E^{ego}_{T}$ severs as query, $E^{ego}_{V}$ works as key and value. The fused result $CA_{fuse}$ is given by:
$\text{CA}_{\text{fuse}}=\textit{CrossAtt}(E^{ego}_{\text{T}}W_{Q},E^{ego}_{\text{V}}W_{K},E^{ego}_{\text{V}}W_{V})$ (1)
where CrossAtt() means the cross attention, $W_{Q}$, $W_{K}$, $W_{V}$ denote the learnable parameters. After that, we combine the $CA_{fuse}$ with the initial $E^{ego}_{V}$ in a residual connection way with the learnable weight $k_{lea}$ as,
$E^{ego}_{MM}=k_{lea}\cdot E^{ego}_{V}+(1-k_{lea})\cdot CA_{fuse}$ (2)
We highlight that though the base PSALM enables various condition prompts, they are never being used at the same time. While our MCFuse module proposes first to generate the text description for the masked object and then manage to fuse the multimodal conditions together.
3.3 Key Module: Cross-View Object Alignment
The Cross-View Object Alignment (XObjAlign) is mainly proposed to enhance the model’s consistent understanding of the same object but from totally different views i.e., ego-centric and exo-centric. The main insight behind this module is that a robust model should have a similar embedding for ego-masked object and also exo-masked object, and also considering the fact that the ego visual emb $E^{ego}_{V}$ is working as the condition that directly affect the prediction of the Mask Generator. Thus, we propose to apply the consistency constraint exactly in the object visual embedding space.
As stated in the overview, to achieve such a consistency constrain, we propose to extract both the ego visual emb $E^{ego}_{V}$ and exo visual emb $E^{exo}_{V}$, after that, the only thing left for the XObjAlign is to compare the similarities of these two embeddings and resulting in the cross-view object consistency loss $\mathcal{L}_{Xobj}$ as,
$\mathcal{L}_{Xobj}=\textit{Sim}(E^{ego}_{V},E^{exo}_{V})$ (3)
where the Sim() represents the similarity of the embedding features, and we use Euclidean distance to calculate it.
We highlight that though its simple design, the idea of aligning the cross-view object is worthy and our XObjAlign module enhances the model’s robustness on the view shift issue with no single additional parameter introduced.
@@ 4 Experiments
Datasets. We validate our method on the large-scale Ego-Exo4D [10], using its benchmark for ego-exo object correspondence tasks. The dataset contains 1.8 million annotated object masks at 1 fps across 1335 video takes, covering domains like Cooking, BikeRepair, Health, Music, Basketball, and Soccer. Each paired video features an average of 5.5 objects, tracked for an average of 173 frames. We use the standard Train/Val/Test splits but make the following adjustments: 1) To focus on cross-view segmentation, we retain only data where objects appear in both views; 2) We introduce a “Small” TrainSet (about one-third of the original “Full” TrainSet) for improved efficiency and data storage; 3) Testing is conducted on the Val set due to the lack of ground truth for the Test set. Our processed splits will be made available for future comparisons.
Network Modules. Following PSALM [68], we use Swin-B [37] as the Visual Encoder and Phi-1.5 1.3B [31] as the LLM. The MLP Projector, consisting of Conv2d, BatchNorm2d, and FC layers, maps the visual latent space to the LLM space. The Pixel Decoder and Mask Generator are adapted from Mask2Former [1]. Mask Token Extraction is the same as PSALM, mainly applying average pooing of mask upon image features. MCFuse is described in Sec. 3.2, and XObjAlign requires no parameters.
Implementation Details. The pre-trained PSALM model serves as our initialization. In S1 training, we use /${}^{1}\!/\!_{20}$ of the training data to train MCFuse. In S2 training, the full training set is used for optimizing all modules except the Swin-B vision encoder. Training details such as epochs, batch size, and learning rate are provided in the Appendix. All training runs on 4xA100 GPUs, while testing is conducted on a single A100, A6000, or L4 GPU. Training on Small TrainSet takes 5-6 hours, while training on the Full TrainSet takes around 20 hours.
@@
Method ZSL Type TrainSet IoU $\uparrow$ Method ZSL Type TrainSet IoU $\uparrow$
XSegTx [10] S - 0.3 XSegTx [10] S - 1.3
XSegTx [10] S Full 6.2 XSegTx [10] S Full 30.2
XView-Xmem [10] ST - 16.2 XView-Xmem [10] ST - 13.5
XView-Xmem [10] ST Full 17.7 XView-Xmem [10] ST Full 20.7
XView-Xmem + XSegTx [10] ST Full 36.9 XView-Xmem + XSegTx [10] ST Full 36.1
SEEM [69] S - 1.1 SEEM [69] S - 4.1
PSALM [68] S - 7.9 PSALM [68] S - 9.6
PSALM [68] S Small 39.7 PSALM [68] S Small 44.1
PSALM [68] S Full 41.3 PSALM [68] S Full 47.3
ObjectRelator (Ours) S Small 44.3 ObjectRelator (Ours) S Small 49.2
ObjectRelator (Ours) S Full 43.9 ObjectRelator (Ours) S Full 50.9
Table 2: Comparison results on Ego-Exo4D object correspondence benchmark. All the results are tested on the Val set. $\circ$ means results from PSALM [68], $\bullet$ means results are reported by us.
@@ 4.1 Main Results on Ego-Exo4D
Baselines and Competitors. Directly aligned prior methods are quite limited, while still some efforts have been made. Specifically, 1) Ego-Exo4D [10] forms a spatial-based “XSegTx” method by adapting SegSwap [51], and two spatiotemporal-based “XView-XMem”, “XView-XMem + XSegTx” via adapting XMem [3] and combing it with “XSegTx”; 2) PSALM [68] tests its ZSL results; 3) In this paper, we also adapt and test the ZSL capabilities of SEEM [69], one additional universal segmentation method that decodes visual prompts by referring to the image. The adaption is mainly made by using the query view (e.g., ego in Ego2Exo) to generate the prompt for the target view (e.g., exo in Ego2Exo). Additionally, we retrain PSALM to form a more competitive baseline.
@@
Results on Ego-Exo4D. The comparison results are summarized in Tab. 2, we indicate whether the method employs ZSL learning, whether it is spatial-based (S) or spatio-temporal (ST) based, the TrainSet used, and most importantly its IoU metric. From the results, we first notice that our proposed ObjectRelator model significantly improves the base PSALM and achieves state-of-the-art performance on the ego-exo object correspondence task. Take models trained on Small TrainSet as an example, our ObjectRelator improves the PSALM from 39.7 to 44.3 on Ego2Exo and from 44.1 to 49.2 on Exo2Ego. These results well demonstrate the effectiveness of our proposed approach.
@@
In addition to our SOTA results, several other key observations are as follows: 1) ZSL Inference Performance: Results show that nearly all the tested methods, including universal segmentation models like SEEM and PSALM, struggle to generalize smoothly to the ego-exo object correspondence task. While SEEM and PSALM excel on other out-of-domain tasks, such as open-vocabulary segmentation [6] and video object tracking [42], their performances are degraded here, highlighting the significant challenges posed by view shifts between ego and exo videos. This underscores that simply applying existing models, even those extensively pre-trained ones, is insufficient. Dedicated models specifically designed and trained for cross-view scenarios remain essential. XView-XMem shows the best ZSL performance, likely due to its video object tracking foundation. However, this advantage diminishes in the retrained XView-XMem. 2) Baseline Retraining: After retraining models on the benchmark, all baselines show reasonable performance, further emphasizing the importance of model adaptation for this task. In general, the superior performance of XView-Xmem and “XView-Xmem + XSegTx” over XSegTx suggests that temporal information aids cross-view alignment. However, this advantage can be offset by the model designs, as demonstrated by PSALM. The success of PSALM shows the effectiveness of interactive prompting, making it a strong baseline choice. 3) Differences in Query Types: Across models, performance is in general better on Exo2Ego than Ego2Exo, indicating that tasks with an ego-view query are more challenging. XSegTx exemplifies this disparity, achieving an IoU of 30.2 on Exo2Ego but only 6.2 on Ego2Exo. This aligns with intuitive expectations, as exo-view objects tend to be smaller and embedded within complex environments, increasing segmentation difficulty, while ego-view objects appear larger and more distinguishable from the background. 4) Impact of Train Set Size: Models trained on the Full TrainSet generally perform better, but we also observe that the Small TrainSet can effectively serve as a testbed for evaluating approaches, with results that approach and even occasionally exceed those of the Full TrainSet. This demonstrates the rationality of using Small TrainSet for efficient experimentation and evaluation, making it a practical alternative for benchmarking cross-view methods while reducing data requirements.
@@ 4.2 Ablation on Proposed Modules
Effectiveness of the Proposed Modules. To evaluate the contributions of every key component. Starting with the base PSALM, we incrementally added the MCFuse and XObjAlign to observe their individual impact. Results are reported in Tab. 3, in addition to the IoUs, the numbers of model parameters (No.Param.) are also provided.
@@
Method MCFuse XObjAlign Ego2Exo $\uparrow$ Exo2Ego $\uparrow$ No.Param.
@@
Further Ablations on Module Options. To thoroughly assess the optimality of our design, we conduct additional ablation experiments applying various options to our modules. These include testing: whether cross attention (CA) is the best method for fusing conditions; the impact of adopting a learnable residual connection; the usage of the learnable weight; and the most suitable placement for applying XObjAlign. The results are presented in Tab. 4.
For the fusion-related experiments, we designed four variants: “Add”, “CA w/o Params”, “CA, S2”, “CA + LearnResidual, S2”. The “Add” method simply adds the ego text and visual conditions as an easy test of our generated text descriptions and also the idea of multi-condition fusion. The “CA w/o Params” variant applies cross-attention without parameters, formulated as $(E^{ego}_{T}\cdot{E^{ego}_{V}}^{T})\cdot E^{ego}_{V}$, which helps evaluate model sensitivity to added parameters. “CA, S2” represents standard cross-attention fusion, as shown in Eq. 1. “CA + LearnResidual, S2” uses our MCFuse module. The “S2” denotes that the entire network is trained only once (i.e., our second training stage). Results show: 1) Even simple methods like “Add” and “CA w/o Params” improve upon the base model, validating MCFuse’s core concept. 2) The notable drop with “CA, S2” suggests caution when introducing new parameters into other pre-trained modules, motivating our two-stage training strategy. This strategy is further validated by comparing “CA + LearnResidual, S2” with our MCFuse. 3) The improvement from “CA, S2” to “CA + LearnResidual, S2” supports protecting the more reliable visual prompt condition. For the residual weight experiments, we compare our learnable approach with fixed weights $k$ (0.2, 0.5, 0.8). Results indicate that learnable weight $k_{lea}$ reduces manual tuning while effectively adapting to appropriate values. Regarding the placement of XObjAlign, we compare put XObjAlign “After MCFuse” with “Before MCFuse”. Results show the latter is more effective, likely because applying the alignment loss $\mathcal{L}_{XObj}$ upon MCFuse may skew optimization toward better fusion objectives.
Ablation Factor Method IoU $\uparrow$
PSALM [68] 39.7
Fusion Add 42.6
CA w/o Params 42.1
CA, S2 22.0
CA+LearnResidual, S2 41.3
CA+LearnResidual, S1-2 (Ours) 43.2
Residual Weight $k=0.2$ 42.1
$k=0.5$ 41.7
$k=0.8$ 43.1
$k_{lea}$ (Ours) 43.2
Placement of XObjAlign After MCFuse 43.6
Before MCFuse (Ours) 44.3
Table 4: Ablation study of proposed modules. Models are trained on Ego2Exo Small TrainSet.
4.3 More Analysis
@@
Table 5: Results of Joint Training on Small TrainSets.
@@
Method Training Testing Conditions Ego2Exo Testing Exo2Ego Testing

### v2

@@ 3.2 Multimodal Condition Fusion
Our Multimodal Condition Fusion (MCFuse) module is mainly presented to incorporate text descriptions into the model, serving as an additional prompt alongside visual cues to enhance the model’s ability on locating target objects. There are mainly two considerations for the design of MCFuse, 1) Given the effectiveness of cross-attention in capturing intricate relationships across modalities, it serves as an ideal choice for fusing our multimodal conditions; 2) the model-generated texts can sometimes be unreliable, prioritizing visual information is prudent. This motivates our use of a residual design, with the visual input serving as the primary pathway. Additionally, to eliminate the need for manual tuning of the residual pathway’s weight, we introduce a learnable parameter, $k_{lea}$.
@@
Formally, as shown in Fig. 3, given the ego text emb $E_{txt}^{*}\in 1\times D$ and ego visual emb $E_{vis}^{*}\in N\times D$ as input, the cross-attention is applied to the broadcasted $E_{txt}^{*}$ and $E_{vis}^{*}$, with $E_{txt}^{*}$ severs as query, $E_{vis}^{*}$ works as key and value. The fused result $CA_{fuse}$ is given by:
$\textit{CA}_{\text{fuse}}=\operatorname{\textit{CrossAtt}}(E_{txt}^{*}W_{Q},E_{vis}^{*}W_{K},E_{vis}^{*}W_{V}),$ (4)
where CrossAtt() means the cross attention, $W_{Q}$, $W_{K}$, $W_{V}$ denote the learnable parameters. After that, we combine the $CA_{fuse}$ with the initial $E_{vis}^{*}$ in a residual connection way with the learnable weight $k_{lea}$ as,
$E_{con}^{*}=k_{lea}\cdot E_{vis}^{*}+(1-k_{lea})\cdot CA_{fuse}.$ (5)
Although the cross-attention is not new, we highlight that: 1) We are the first to introduce text modality into this task and leverage it to enhance object localization. Results in Tab. 3 and Fig. 5 well show its effectiveness. 2) While prior segmentation models, including the base PSALM, allow various condition prompts, they do not utilize them simultaneously. In contrast, our approach integrates multimodal prompts to refer to the same object(s).
3.3 Cross-View Object Alignment
The Cross-View Object Alignment (XObjAlign) is mainly proposed to enhance the model’s consistency across totally different views. The main insight behind this module is that a robust model should generate similar embeddings for both ego-masked and exo-masked object(s). Furthermore, since the ego visual embedding $E_{vis}^{*}$ serves as a conditioning factor, its quality directly affects the predictions of the Mask Generator. Thus, we propose to apply the consistency constraint exactly in the object-level embedding space.
As in Sec. 3.1, to achieve such an consistency constraint, we extract both the ego visual emb $E_{vis}^{*}$ and exo visual emb $E_{vis}$, after that, the only thing left for XObjAlign is to compare the similarities of these two embeddings and resulting in the cross-view object consistency loss $\mathcal{L}_{Xobj}$ as,
$\mathcal{L}_{Xobj}=\operatorname{\textit{Dist}}(E_{vis}^{*},E_{vis}),$ (6)
where the Dist() represents the Euclidean distance. Without introducing additional parameters or supervision, our XObjAlign module ensures cross-view consistency, effectively handling object alignment across different viewpoints.
@@ 4 Experiments
Datasets. We validate our method mainly on the large-scale Ego-Exo4D [13], using its benchmark for ego-exo object correspondence tasks. The dataset contains 1.8 million annotated object masks at 1 fps across 1335 video takes, covering domains like Cooking, BikeRepair, Health, Music, Basketball, and Soccer. Each paired video features an average of 5.5 objects, tracked for an average of 173 frames. We use the standard Train/Val/Test splits but make the following adjustments: 1) To focus on cross-view segmentation, we retain only data where objects appear in both views; 2) We introduce a “Small” TrainSet (about one-third of the original “Full” TrainSet) for improved efficiency and data storage; 3) Testing is conducted on the Val set due to the lack of ground truth for the Test set. Our processed splits will be made available for future comparisons. To evaluate methods more comprehensively, we introduce HANDAL-X, another cross-view object segmentation benchmark adapted from the HANDAL [14]. More details are in the Supplementary Materials (Supp. Mat.)
Network Modules. Following PSALM [75], we use Swin-B [42] as the Visual Encoder and Phi-1.5 1.3B [36] as the LLM. The MM Projector, consisting of Conv2d, BatchNorm2d, and FC layers, maps the visual latent space to the LLM space. The Pixel Decoder and Mask Generator are adapted from Mask2Former [2]. MCFuse is described in Sec. 3.2, and XObjAlign requires no parameters.
Metrics. Following standard segmentation tasks, we use IoU as the primary metric. We also report Location Error (LE), Contour Accuracy (CA), and Visibility Accuracy (VA) for main results on Ego-Exo4D as proposed in [13]. Briefly, LE measures the normalized distance between centroids of predicted and ground-truth masks, CA evaluates mask shape similarity, and VA assesses model’s ability to estimate object existence visibility in the target view.
Implementation Details. The pre-trained PSALM model serves as our initialization. In S1 training, we use /${}^{1}\!/\!_{20}$ of the training data to train MCFuse. In S2 training, the full training set is used for optimizing all modules except the Swin-B vision encoder. Training details such as gpus, epochs, batch size, and learning rate are provided in the Supp. Mat.
@@
[…]
XSegTx [13] - 0.3 0.149 0.012 98.9 XSegTx [13] - 1.3 0.188 0.023 98.0
XSegTx [13] Full 6.2 0.091 0.171 96.5 XSegTx [13] Full 30.2 0.107 0.391 96.9
XView-Xmem [13] - 16.2 0.145 0.212 61.9 XView-Xmem [13] - 13.5 0.153 0.178 34.4
XView-Xmem [13] Full 17.7 0.146 0.259 85.0 XView-Xmem [13] Full 20.7 0.140 0.293 75.2
Xmem+XSegTx [13] Full 36.9 0.068 0.447 85.6 Xmem+XSegTx [13] Full 36.1 0.117 0.438 96.0
SEEM [77] - 1.1 0.146 0.052 99.3 SEEM [77] - 4.1 0.183 0.078 99.2
PSALM [75] - 7.9 0.156 0.178 99.3 PSALM [75] - 9.6 0.140 0.177 99.3
PSALM [75] Small 39.7 0.058 0.537 99.8 PSALM [75] Small 44.1 0.067 0.553 99.7
PSALM [75] Full 41.3 0.065 0.536 99.8 PSALM [75] Full 47.3 0.067 0.571 99.7
ObjectRelator (Ours) Small 44.3 0.051 0.546 99.9 ObjectRelator (Ours) Small 49.2 0.065 0.577 99.8
ObjectRelator (Ours) Full 45.4 0.056 0.548 99.9 ObjectRelator (Ours) Full 50.9 0.064 0.586 99.8
Table 2: Comparison results on Ego-Exo4D object correspondence benchmark. All the results are tested on the Val set. $\circ$ means results from PSALM [75], $\bullet$ means results are reported by us. We use “Xmem+XSegTx” for short of “XView-Xmem+XSegTx”.
@@ 4.1 Main Results on Ego-Exo4D
Baselines and Competitors. Directly aligned prior methods are quite limited, while still some efforts have been made. Specifically, 1) Ego-Exo4D [13] forms a spatial-based “XSegTx” method by adapting SegSwap [57], and two spatiotemporal-based “XView-XMem”, “XView-XMem+XSegTx” via adapting XMem [4] and combing it with “XSegTx”; 2) PSALM [75] tests its ZSL results; 3) In this paper, we also adapt the SEEM [77], one additional universal segmentation method that decodes visual prompts by referring to the image. The adaptation is mainly made by using the query view (e.g., ego in Ego2Exo) to generate the prompt for the target view (e.g., exo in Ego2Exo). 4) Additionally, PSALM is retrained as a competitive baseline.
@@
Results on Ego-Exo4D. The comparison results are summarized in Tab. 2. We indicate whether the method employs ZSL learning, the TrainSet used, and several metrics: IoU, LE, CA, and VA. From the results, we first notice that our proposed ObjectRelator model significantly improves the base PSALM and achieves state-of-the-art performance on the ego-exo object correspondence task. Take models trained on Small TrainSet as an example, our ObjectRelator improves the PSALM from 39.7 to 44.3 on Ego2Exo and from 44.1 to 49.2 on Exo2Ego. Regarding LE, CA, and VA, Our ObjectRelator consistently outperforms PSALM, achieving clear margins over all other competitors. These results well demonstrate the effectiveness of our approach.
@@
In addition to our SOTA results, several other key observations include: 1) Different Metrics: Although IoU, LE, and CA are defined differently, they all fundamentally measure the discrepancy between the predicted and ground truth masks. Our results also show high consistency among them. Additionally, most of the models exceed 96% on VA, indicating that object visibility estimation is relatively easy. Thus, we mainly use IoU for further analysis. 2) ZSL Inference Performance: Results show that nearly all the tested methods, including universal segmentation models like SEEM and PSALM, struggle to generalize smoothly to this task. While SEEM and PSALM excel on other out-of-domain tasks, such as open-vocabulary segmentation [8] and video object tracking [47], their performances are degraded here. This underscores that simply applying existing models, even those extensively pre-trained ones, is insufficient. Dedicated models specifically designed and trained for cross-view scenarios remain essential. XView-XMem shows the best ZSL performance, likely due to its video object tracking foundation. However, this advantage diminishes in the retrained XView-XMem. 3) Baseline Retraining: After retraining models on the benchmark, all baselines show reasonable performance, further emphasizing the importance of model adaptation for this task. In general, the superior performance of XView-Xmem and “Xmem+XSegTx” over XSegTx suggests that temporal information aids cross-view alignment. However, this advantage can be offset by the model designs, as demonstrated by PSALM. The success of PSALM makes it a strong reason for choosing it as our baseline. 4) Differences in Query Types: Across models, performance is in general better on Exo2Ego than Ego2Exo, indicating that ego-view queries are more challenging. XSegTx exemplifies this disparity, achieving an IoU of 30.2 on Exo2Ego but only 6.2 on Ego2Exo. This aligns with intuitive expectations, as exo-view objects tend to be smaller and embedded within complex environments, increasing segmentation difficulty, while ego-view objects appear larger and more distinguishable. 5) Impact of Train Set Size: Models trained on the Full TrainSet generally perform better, but the Small TrainSet also yields comparable results. This demonstrates that the Small TrainSet is a practical alternative for benchmarking this task while reducing data requirements.
@@ 4.2 Ablation on Proposed Modules
To evaluate the contribution of each component, starting with the base PSALM, we incrementally add the MCFuse and XObjAlign to observe their individual impact. Results are reported in Tab. 3, in addition to the IoUs, the numbers of model parameters (No.Param.) are also provided.
@@
[…]
More ablations, including multimodal fusion strategies, the impact of the learnable residual connection, comparisons of learnable $k_{lea}$ vs. fixed ratio, and different loss weightings between $\mathcal{L}_{mask}$, $\mathcal{L}_{X{obj}}$, are in the Supp. Mat.
4.3 More Analysis on Ego-Exo4D
@@
Table 4: Results of Joint Training on Small and Full TrainSets.
@@
[…]
@@
4.4 Main Results on HANDAL-X
To further validate our method, we conduct experiments on HANDAL-X. The comparison results, summarized in Tab. 6, include XSegTx, SEEM, and PSALM as competitors. Note that ”XView-XMem” and ”XMem+XSegTx” are excluded since HANDAL-X is a cross-view image segmentation dataset. To differentiate models using different checkpoints, we also indicate the primary datasets used.
[…]
ZSL XSegTx [13] Self-Generated Pairs from COCO 1.5
SEEM [77] COCO Panoptic, RefCOCO/+/g 2.5
PSALM [75] COCO Panoptic, RefCOCO/+/g, … 14.2
PSALM [75] Ego-Exo4D 39.9
ObjectRelator Ego-Exo4D 42.8
Retrained PSALM [75] HANDAL-X 83.4
ObjectRelator HANDAL-X 84.7
Table 6: Comparison results on HANDAL-X benchmark. $\circ$ means checkpoints from priors works; $\bullet$ means models trained in this paper; For the PSALM and ObjectRelator trained on Ego-Exo4D, the checkpoints with Exo2Ego Small TrainSets are used.
We observe that: 1) This benchmark is generally easier than Ego-Exo4D, where extreme viewpoint changes are considered. However, it also serves as a valuable testbed for more common scenarios. 2) ZSL results follow a similar trend to Ego-Exo4D, with PSALM outperforming others, followed by SEEM. 3) A key finding is that models trained on Ego-Exo4D (PSALM and ObjectRelator) demonstrate strong ZSL ability, significantly surpassing those trained on other datasets (e.g., COCO) with an IoU of around 40. This suggests that training on cross-view data improves generalization to new cross-view settings, even with different scenarios and categories. 4) Across both ZSL and retrained settings, ObjectRelator consistently outperforms PSALM. This further validates our approach.
We highlight that although HANDAL-X is relatively simple, it remains valuable: 1) it effectively reveals performance differences (ranging from 1.5 to 84.7); 2) it serves as a testbed for zero-shot learning; and 3) its natural characteristics—robotics-ready, manipulable objects (e.g., a spoon with appropriate size and shape)—make it particularly suitable for robotics-related applications.